Please show calculation to the problem. A auditor reviewed 25 oral surgery insurance claimed from particular surgical office, determining that the mean out-of-pocket patient billing above the rembursed amount was $ 275.66 with a standard deviation of $ 78.11. (a) At the 5 percernt level of significance, does this sample prove a violation of the guideline that the average patient should pay no more than $ 250 out-of-pocket: State your hypothesis and decision rule. (b) Is this a close decision?
Answer by stanbon(75887)

Please show calculation to the problem. A auditor reviewed 25 oral surgery insurance claimed from particular surgical office, determining that the mean out-of-pocket patient billing above the rembursed amount was $ 275.66 with a standard deviation of $ 78.11.
(a) At the 5 percernt level of significance, does this sample prove a violation of the guideline that the average patient should pay no more than $ 250 out-of-pocket: State your hypothesis and decision rule.
Ho: u = 250
H1: u > 250
---
invT(0.95,24) = 1.711
At the 5% level for a one-tail test, reject Ho if t > 1.711
(b) Is this a close decision?
t(275.66) = (275.66-250)/[78.11/sqrt(25)] = 25.66*5/78.11 = 1.64
----
Conclusion: Fail to reject Ho.
Not really a close decision.
Cheers,
Stan H.